OpenStudy (anonymous):

The high school marching band has a total of 100 students. There are 3 times as many girls in the band as boys. How many girls are in the band?

OpenStudy (anonymous):

60 maybe

OpenStudy (anonymous):

Write an equation, with x = no. of boys. Since there are 3times as many girls, there should be 4x total people in the band. 4x = 100 Solve for x to get the number of boys. Then multiply that by 3 to get the number of girls.

OpenStudy (anonymous):

Let b and g be the number of boys and girls respectively. Solve the following for b and g.\[\left\{b+g=100,\frac{g}{b}=3\right\} \]\[\{b=25,g=75\} \]